OpenStudy (anonymous):

Solve the proportion. x-8/5=2/4 A.) 18 B) 5/2 C) 21/2 D.) 9/2

OpenStudy (anonymous):

\[\frac{x-8}{5}=\frac{1}{2}\iff 2(x-8)=5\times 1\]

OpenStudy (anonymous):

IT C right? @satellite73

OpenStudy (anonymous):

yes